2^{nd} EUROPEAN CONGRESS OF MATHEMATICS July 21 – 27, 1996, Budapest, Hungary

PRELIMINARY REGISTRATION FORM

To: J. Bolyai Mathematical Society / ECM2 Fő utca 68, Budapest, Hungary H-1027 Tel: (+36-1) 201-7656, tel/fax: 201-6974 e-mail: ecm2jbms@math-inst.hu

Please send me further information on the 2^{nd} European Congress of Mathematics.

Name (please print): . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Mailing address: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .
. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

E–mail address: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

Please distribute copies of this form among your colleagues. Please return this form to J. Bolyai Mathematical Society as soon as possible, preferably before July 31, 1995. (You may respond by e-mail.)